OpenStudy (anonymous):

evaluate f(x)=7-x/-5+x when f(a+1)

OpenStudy (blurbendy):

f(a+1) = 7 -(a + 1) / -5 + (a +1) = 7 -a -1 / -5 + a + 1 = 6 - a / -4 + a